OpenStudy (anonymous):

What is the x-coordinate of the vertex of y = |4x + 12| + 3 ? A. -3 B. zero C. 3 D. 4

OpenStudy (anonymous):

yo bro, what is the standard form?

OpenStudy (anonymous):

um...@AyedenCarpenter..this is geo?

OpenStudy (bookworm14):

Vertex is where the sign of "inside expression" inside the absolute brackets changes sign, that is passes through zero. Itmeans the vertex is where 4x+12=0 found this on a different site , Hope this Helps - Bree ;)

OpenStudy (bookworm14):

the answer is C.3

OpenStudy (anonymous):

@Dmb1215

OpenStudy (anonymous):

Nope

OpenStudy (dmb1215):

C

OpenStudy (anonymous):

if 4x+12 = 0 then x = -3 It's A

OpenStudy (bookworm14):

Idk, I tried :(

OpenStudy (anonymous):

you were right when 3x+12=0. Just solved for x wrong

OpenStudy (bookworm14):

sorry haven't covered this in school yet :P

OpenStudy (bookworm14):

ohh okay , :)

OpenStudy (bookworm14):

sorry if I confused you @AydenCarpenter14